Sam rows at a speed of 6.5 miles per hour in still water. If Sam rows 9 miles up-stream and then back again in a total of 3.25 hours, what is the speed of the current, in miles per hour?
Answer by jorel1380(3719)

9/(6.5+n)+9/(6.5-n)=3.25
9(6.5-n)+9(6.5+n)=3.25(42.25-n2)
58.5-9n+58.5+9n=137.3125-3.25n2
3.25n2-20.3125=0
32500n2-203125=0
4n2-25=0
(2n+5)(2n-5)=0
n=-5/2 or 5/2
Throwing out the negative answer, we get the speed of the current to be 5/2 mph..